OpenStudy (anonymous):

need help!! -7(x-4)+8x=12+14 x=? I also have a few more ? if u can help

OpenStudy (ahsome):

Equation: \(-7(x-4)+8x=12+14\) Simplify RHS\[-7(x-4)=26\] Expand LHS\[(-7*x)+(-7*-4)=26\]\[-7x+28=26\] Move \(28\) to the RHS\[-7x=26-28\]\[-7x=-2\] Divide both sides by -7\[\frac{-7x}{-7}=\frac{-2}{-7}\]\[x=0.28571...\]

OpenStudy (ahsome):

Assuming first equation was right

OpenStudy (anonymous):

could u help with a few more?

OpenStudy (ahsome):

I am kinda busy, but there are heaps of others who could help you. Just ask. I will redirect some. @iGreen @ganeshie8 @texaschic101 we need your help Sorry I couldn't be more of a service to you, @hello123321123 :(

OpenStudy (itsbribro):

ill try to help!!

OpenStudy (itsbribro):

@hello123321123

OpenStudy (ahsome):

YAY :D

OpenStudy (itsbribro):

do you have any more questions?

OpenStudy (anonymous):

yes, -6.8+2.6+x=9.7 x=?

OpenStudy (anonymous):

can u help?

OpenStudy (itsbribro):

-6.8+2.6+x=9.7 =139/10 x = 13.9

OpenStudy (anonymous):

I messaged u @itsbribro

OpenStudy (itsbribro):

i didnt get a message.

OpenStudy (itsbribro):

are you ok??

OpenStudy (anonymous):

yes @itsbribro I just messaged y'all my ?